OpenStudy (anonymous):

add. simplify if possible 4t/t^2-9 + t/ t-3

OpenStudy (anonymous):

\[4t/(t^2-9) + t/(t-3) = 4/((t+3)(t-3)) + t/(t-3)\]\[(4t+t^2+3t)/(t^2-9) = (t^2+7t)/(t^2-9)\]